OpenStudy (anonymous):

find the remainder when f (x) = x^3 - 14x^2 + 51x -22 is divided by x - 7.

ganeshie8 (ganeshie8):

use synthetic division http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

i know what to do i just can't figure it out

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

or you can use the remainder theorem the remainder of f(x)/(x - p) is the value f(p) ------------------------------------------------------- in other words, if r is the remainder, then r = f(p)

OpenStudy (anonymous):

walkthrough?

jimthompson5910 (jim_thompson5910):

x - p x - 7 so, p = ???

OpenStudy (anonymous):

p = 7?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

plug that into the function

jimthompson5910 (jim_thompson5910):

plug it into f (x) = x^3 - 14x^2 + 51x -22

OpenStudy (anonymous):

hmmm.

OpenStudy (arbershabani97):